[Yong-Xiang Yang, Chia-Hsuan Hsiung, Gui Chin Liu (from Tamkang university), Chia-Jui Chou, Yuzurihara]
Transient behavior by BPC feedback control
Sometimes, we observed the transient behavior in the error and feedback signal of BPC for ITMX/ETMX/ITMY/ETMY at the same time.
The error signal is computed from the K1:LSC-DARM_IN1_DQ by demodulation of 865Hz for ITMX and ETMX, 845 Hz for ITMY and ETMY, and 870 Hz. If some additional noise was induced in K1:LSC-DARM_IN1_DQ, it’s possible to mimic the shift of the beam spot and the feedback signal will be sent to the suspension and makes the drift to somewhere, even though the actual beam spot was not moved. That's the scenario we are considering.
The timeseries of error and feedback signal for the BPC are attached. You can see the transient behabior before the lockloss.
The whitened spectrogram is also attached. Before the locklos, the noise level seems to be increased due to some reason.
Based on the chat with Ushiba-san, the amount of feedback signal to ITMX/ETMX/ITMY/ETMY looks small to move the mirror drastically and make the lockloss. On ther other hand, the feedback to PRM is in the unit of urad. It's pottential to move the mirror a lot and make the lockloss.
- 2024-12-29 22:58:28 UTC (Fig): Whitened Spectrogram around the BPC frequency (Fig)
- 2024-12-29 22:39:37 UTC (Fig): Whitened Spectrogram around the BPC frequency (Fig)
- 2024-12-29 19:41:26 UTC (Fig): Whitened Spectrogram around the BPC frequency (Fig)
- 2024-12-29 19:04:37 UTC (Fig)
- 2024-12-29 17:26:56 UTC (Fig)
- 2024-12-28 20:36:31 UTC (Fig)
- 2024-12-28 18:29:54 UTC (Fig)
Next step
- Check the behavior of POP90 before the lockloss and check if the POP90 decreased by PRM feedback or not.
- Check the spectrum of all frequency range of K1:LSC-DARM_IN1_DQ by changing the time
- Check more lockloss time about whether this phenomena was happening or not